[Hallc_running] [New Logentry] RC Update

brads at jlab.org brads at jlab.org
Mon Feb 4 08:55:03 EST 2019

Logentry Text:
MCC is a little behind last week's whiteboard predictions.  The cryocyle of 1L20 was completed Sunday evening and they started attempting to spin up beam during Owl.  They ran into a number of issues, but no show stoppers yet.

Hall Lockup has been pushed to late Thursday DAY shift, but things will need to go well to hit that mark.  I'll talk to the Target group et al. about pushing the cool down date back and will let any affected TOs know via email.

The current plan is to setup the machine with imbalanced linacs: 1040 in the North, and 1060 in the South.

1L07 has improved gradient and trip rates are OK.

1L20 is still an open question.  There was little progress bringing it online last night, and it is currently bypassed while they continue work on the rest of the machine.  Ops identified a 'commissioning card' was still installed on that module that needs to be swapped out, so it is not yet a lost cause.  MCC reports that if 1L07 holds up (and no other problems arise) the north linac can reach the gradient needed for the program with some headroom.

There is also an excess ~600W of cryo load in the north as of this weekend that Ops is investigating.  Hopefully just a bad insulating vacuum due to a tripped pump or something...


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3647565
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190204/209c1c21/attachment.html>

More information about the Hallc_running mailing list